OpenStudy (anonymous):

Does the series [absolute value of (sin n)]/(n^2+1) converge or diverge?

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \frac{ \left| \sin n \right| }{ n ^{2} + 1 }\]

OpenStudy (xapproachesinfinity):

i forgot how to do this things i guess you do this by comparison test

OpenStudy (xapproachesinfinity):

just a try not sure \[|\sin n|\leq1 \Longrightarrow \frac{|\sin n|}{n^2+1}\leq \frac{1}{n^2+1}\] take the limit as n goes to oo \[\lim_{n\to\infty}\frac{1}{n^2+1}=0\] hence \[\lim_{n\to\infty}\frac{|\sin n|}{n^2+1}=0\] as a result the sum converges

OpenStudy (xapproachesinfinity):

but i'm not sure this correct lol :)

OpenStudy (xapproachesinfinity):

did you do comparison test?

OpenStudy (anonymous):

Thanks!! :)

OpenStudy (anonymous):

I was thinking of the comparison test, but I wasn't quite sure how to do it

OpenStudy (xapproachesinfinity):

welcome :) check your notes on how is this done and see if that's what i did^^

OpenStudy (xapproachesinfinity):

@surjithayer

OpenStudy (anonymous):

None of the examples from my notes are with absolute values, but the make up series is of the highest power, so I think the regular series should be \[\frac{ 1 }{ n ^{2} }\]

OpenStudy (xapproachesinfinity):

well i know 1/(n^2+1) converges since 1/n^2 does converge as well

OpenStudy (anonymous):

So, you're supposed to leave the absolute value of sine alone?

OpenStudy (xapproachesinfinity):

hmm not sure what you meant by that?

OpenStudy (anonymous):

Uhmm during the comparison test, what are you supposed to do with 'sin n'?

OpenStudy (xapproachesinfinity):

bring help @myininaya

OpenStudy (xapproachesinfinity):

well we are supposed to compare the giving series with a different series that we are familiar with and that's what i did above

OpenStudy (xapproachesinfinity):

if that series that we know converges so is our series if not then the series deos not converge this is what comparison test is about

OpenStudy (anonymous):

So would it start like this: \[\frac{ \left| \sin n \right| }{ n ^{2}+1 } \le \frac{ \left| \sin n \right| }{ n ^{2} }\]

OpenStudy (xapproachesinfinity):

refer to this page, there are some examples of comparison test

OpenStudy (xapproachesinfinity):

no see my work above

OpenStudy (anonymous):

But that isn't quite like what my teacher showed us

OpenStudy (xapproachesinfinity):

well i don't know what your teacher showed you but what i did above is correct! i'm just not sure if i need more steps to say that it is convergent

OpenStudy (xapproachesinfinity):

if you look at the link i posted they are doing the same thing

OpenStudy (anonymous):

I don't see a link

OpenStudy (xapproachesinfinity):

http://tutorial.math.lamar.edu/Classes/CalcII/SeriesCompTest.aspx

OpenStudy (xapproachesinfinity):

i thought i posted it

OpenStudy (xapproachesinfinity):

haha xD

OpenStudy (xapproachesinfinity):

\[\large \frac{|\sin n|}{n^2}\] how would you prove this convergent or not

OpenStudy (xapproachesinfinity):

that's what your were thought in class?

OpenStudy (anonymous):

Well, we were only taught with numerators of 1 and that's why I'm not sure what to do with

OpenStudy (anonymous):

to do with sin n *

OpenStudy (xapproachesinfinity):

well look above i have 1/ (n^2+1) and that's what you needed this series converge just line 1/n^2

OpenStudy (anonymous):

But what about \[\left| \sin n \right|\] ?

OpenStudy (xapproachesinfinity):

hehe okay you didn't got what i did by definition of sin we have: \[|\sin n|\leq1\] yes

OpenStudy (xapproachesinfinity):

?

OpenStudy (xapproachesinfinity):

agree?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

okay then i multiplied the inequality by \[\frac{1}{n^2+1}\] which is positive so no change in the inequaltiy i get \[\large \frac{|\sin n|}{n^2+1}\leq\frac{1}{n^2+1}\] no get the form i needed \[\Large a_n\leq b_n\] since\[\Large b_n\] converges then \[\large a_n\] also converges

OpenStudy (xapproachesinfinity):

now*

OpenStudy (xapproachesinfinity):

does it make sense now?

OpenStudy (anonymous):

Yeahh! thanks!

OpenStudy (xapproachesinfinity):

welcome!